# A Calculus of Inheritance

### v4

@@ Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. It trivially embeds the $\lambda$-calculus, although the entire semantics rests solely on naive set theory; as a consequence, all constructs including inheritance are inherently commutative, idempotent, and associative; the linearization problem of multiple inheritance does not arise. This induces a fully abstract semantics of the lazy $\lambda$-calculus with respect to Böhm tree equivalence (Barendregt, 1984). Inheritance-calculus is distilled from MIXINv2, a practical implementation in which we observed further emergent phenomena: the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); self-reference resolves to multiple targets; and programs are immune to the Expression Problem (Wadler, 1998). This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991). These properties suggest applications to configuration languages, dependency injection, object-oriented programming, composable effect systems, modular software architectures, file-system-as-compiler, general-purpose programming, and no-code development.
declarative programming, inheritance, denotational semantics, self-referential records, expression problem, $\lambda$-calculus, Böhm trees, configuration languages
@@
$c$s: Theory of computation Denotational semanticsccs: Theory of computation Object oriented constructsccs: Theory of computation Functional constructsccs: Theory of computation Program schemesccs: Software and its engineering Formal language definitionsccs: Software and its engineering Extensible languagesccs: Software and its engineering Object oriented languages
@@ 1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging. Among these, the NixOS module system stands out: its recursive attribute set merging with fixed-point semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so expressive.
@@
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures, that is, initial algebras in the sense of universal algebra. We challenge this assumption: configuration values are better understood as lazily observable, possibly infinite structures, where semantics is determined by the observer and a finite prefix suffices for any finite observation. This is not merely a theoretical distinction: the nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
@@
In developing real programs in MIXINv2, including boolean logic and natural number arithmetic (Section 5) and a web application (Section Function color blindness: ), we found that the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) is structurally absent: because records are sets, inheritance is inherently commutative, idempotent, and associative (Section 3). Self-reference naturally resolves to multiple objects rather than one; prior systems assume single-object resolution and reject the multi-path situation (Scala, NixOS module system). Self-referential records require no fixed-point combinator—the inheritance tree is the call stack, and qualified this resolution is a first-order set-theoretic query, not a fixed-point construction. We also found that boolean logic and arithmetic, implemented as separate MIXINv2 files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism; that Church-encoded values automatically form trie structures; and that operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$, which is the relational semantics of logic programming (van Emden and Kowalski, 1976) (Section 5); and that the same web application MIXINv2 file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015), equivalently CPS-agnostic behavior (Section Function color blindness: ).
@@
This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming. Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into inheritance-calculus, inheritance-calculus is Turing complete. The translation reveals an asymmetry: embedding the $\lambda$-calculus into inheritance-calculus requires five local, syntax-directed rules (Section 4), but the reverse is not macro-expressible under the equivalence notions fixed in Section 4.2 and Appendix A— inheritance-calculus enables open extension that changes operational equivalence on $\lambda$-terms, establishing a strict separation in Felleisen’s sense of expressiveness (Felleisen, 1991) (Section 4.2).
@@
All constructions in this paper have been implemented and tested in MIXINv2, available as open source (Yang, 2026), with a test suite covering every example. [^1] MIXINv2 is statically typed and therefore stricter than the untyped inheritance-calculus. All examples in this paper are ported from MIXINv2 and conform to both semantics. Elements added to satisfy the type checker are harmless for the algorithms presented here and may aid readability. All tests are mechanically verified against expected output snapshots.
@@
  • We present inheritance-calculus, a minimal computational model for declarative programming that contains only three constructs (record, definition, inheritance) and no functions or scalar types.
  • Inheritance operates at the tree level: a record is a set of elements, both definitions and inheritance sources, so inheriting from two mixins that share a label recursively merges their subtrees (deep merge), rather than overriding at the method level as in prior mixin and trait calculi. Since records are sets, inheritance is inherently commutative, idempotent, and associative. The linearization problem inherent in prior mixin and trait calculi does not arise.
  • The semantics is purely observational: there are no reduction rules, no evaluation order, and no notion of a result that a term reduces to. The observer drives computation by querying paths in a lazily constructed tree.
  • Self-referential records require no fixed-point combinator. The $\lambda$-calculus translation maps each application to an inheritance; the resulting inheritance tree is the call stack. The $\mathtt{this}$ function of Section 3 resolves self-references by walking this tree: a first-order set-theoretic query, not a fixed-point construction.
  • Self-reference resolves to multiple objects, not one. Path-dependent types with inheritance naturally produce this situation; prior systems assume single-object resolution and reject it (Scala and the NixOS module system, Appendix B). Since inheritance is idempotent, all objects contribute equally; this semantics does not introduce a route-priority disambiguation rule.
  • The $\lambda$-calculus (in A-normal form) translates into inheritance-calculus in five rules (Section 4). The translation is fully abstract with respect to Böhm tree equivalence under $\lambda$-definable contexts (Section 4.1), yielding a semantics of the lazy $\lambda$-calculus that requires no domain theory.
  • We observe further properties: extensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section Function color blindness: ).
  • Open extension via inheritance changes the operational equivalence of $\lambda$-terms, establishing via Felleisen’s Theorem 3.14 (Felleisen, 1991) that the $\lambda$-calculus cannot macro-express inheritance-calculus (Section 4.2).
@@ 2. Syntax
Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a non-negative integer (path length).
@@
The $\mapsto$ in $\ell\mapsto e$ defines a property, not a let-binding. There is no variable binding in inheritance-calculus. MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; scalars did not arise.
@@ 3. Mixin Trees
Given an AST, what are the properties visible at each position in the tree? This section answers the question by defining a collection of mutually recursive functions over paths.
A mixin tree has a single abstraction, the mixin: every mixin can serve as field, method, class, and object at the same time in object-oriented terms.
Unlike the $\lambda$-calculus, which requires $\beta$-reduction to specify how terms compute, inheritance-calculus has no reduction steps: no rewriting of terms, no evaluation order, and no notion of a result that a term reduces to. Instead, the semantics is purely observational: one queries whether a label $\ell$ is a property at a path $p$ in the inheritance tree. The mutually recursive functions defined below, namely $\mathtt{properties}$, $\mathtt{supers}$, $\mathtt{overrides}$, $\mathtt{bases}$, $\mathtt{resolve}$, and $\mathtt{this}$, constitute the entire semantics.
Each function has an analogue in object-oriented languages: $\mathtt{Path}$ identifies a position in the tree, $\mathtt{init}$ and $\mathtt{last}$ navigate paths, $\mathtt{defines}$ extracts the locally defined members from the AST, $\mathtt{inherits}$ extracts the inheritance declarations from the AST, $\mathtt{properties}$ computes all visible members at a path including both locally defined and inherited ones, $\mathtt{supers}$ computes the transitive inheritance closure, $\mathtt{bases}$ collects the direct base classes, and $\mathtt{overrides}$ generalizes method override to the tree setting, collecting the identity-sharing paths that arise when inheritance introduces multiple definitions of the same label at the same scope level. Two functions, the reference resolution function $\mathtt{resolve}$ and the qualified this resolution function $\mathtt{this}$, have object-oriented analogues whose semantics differ significantly; they are discussed in detail below.
All functions are pure and may be cached (memoised) without changing semantics.
@@
A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ell_{2},\ldots,\ell_{n})$ that identifies a position in the tree. [^2] Since paths are referentially transparent, an implementation may intern paths so that structural equality reduces to pointer equality. We write $p$ for a path. The root path is the empty sequence $()$. Given a path $p$ and a label $\ell$, $p\mathbin{\triangleright}\ell$ is the sequence $p$ extended with $\ell$. For a non-root path, the parent path $\mathtt{init}(p)$ is $p$ with its last element removed, and the final label $\mathtt{last}(p)$ is the last element of $p$. For convenience, we sometimes write $\ell$ for $\mathtt{last}(p)$ when the path is clear from context. With paths in hand, we can state what the AST provides at each path.
@@
Parsing populates $\mathtt{defines}$ and $\mathtt{inherits}$ from each element of a record. A definition $\ell\mapsto e$ contributes $\ell$ to $\mathtt{defines}(p)$. A reference contributes $(n,\;\ell_{\mathrm{down},*})$ to $\mathtt{inherits}(p)$. Since records are sets, the order of elements is irrelevant and duplicate elements have no effect. Both $\mathtt{defines}$ and $\mathtt{inherits}$ are pure data; they are not functions of runtime state. Given these two primitives, we can now answer the question posed at the beginning of this section.
@@
Because inheritance-calculus has no separate notion of “method,” a mixin simultaneously plays the role of a class and the role of a method in object-oriented languages. When a mixin is viewed as a method, a child scope should be able to find it through the parent’s label even if the child does not redefine that label—this is method override. Concretely, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overrides}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge. The overrides of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
@@
At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a non-negative integer, the recursion terminates.
@@
Observational semantics without reduction
The definitions above constitute the complete semantics of inheritance-calculus. There are no reduction rules, no evaluation contexts, and no small-step or big-step judgments. In the $\lambda$-calculus, a term such as $(\lambda x.\,x)\;y$ must be reduced to $y$ before its meaning is apparent; the operational semantics specifies how and in what order reductions occur. In inheritance-calculus, an inheritance tree such as $\{a\mapsto\{\},\;b\mapsto\{\}\}$ does not reduce to anything: it is the value. The functions $\mathtt{properties}$, $\mathtt{supers}$, etc. determine what is observable at each path, and that is all. Computation arises not from rewriting terms but from querying an inductively deeper path in a lazily constructed tree: the observer drives the computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand. Formally, the mutually recursive equations (1)–(6) define set membership inductively: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. The corresponding operator on the product lattice of path sets is monotone, and its least fixed point exists by the Knaster–Tarski theorem (Tarski, 1955). The detailed well-definedness argument, which identifies the lattice, verifies positivity of each equation, and handles the transitive closure $\mathtt{bases}^{*}$, is given in Appendix C. When no finite derivation exists for a given query, the membership is not in this least fixed point and the semantics is undefined at that point, analogous to divergence in the $\lambda$-calculus, and the expected consequence of Turing completeness. Identity is determined by observable behavior rather than by internal structure.

### v6

@@ Abstract.
Just as the $\lambda$-calculus uses three primitives (abstraction, application, variable) as the foundation of functional programming, inheritance-calculus uses three primitives (record, definition, inheritance) as the foundation of declarative programming. By unifying classes, methods, objects, and properties under a single record abstraction, the calculus models inheritance simply as set union. Consequently, composition is inherently commutative, idempotent, and associative, structurally eliminating the multiple-inheritance linearization problem. Its semantics is first-order, denotational, and fixpoint-computable. These three properties extend to the lazy $\lambda$-calculus, since Böhm tree equivalence (Barendregt, 1984) is fully abstract for a sublanguage of inheritance-calculus, except that recursive $\lambda$-programs that diverge under $\beta$-reduction may converge under $\mathrm{lfp}(T_{P})$. Inheritance-calculus is distilled from MIXINv2, a practical implementation in which the same code acts as different function colors (Nystrom, 2015); ordinary arithmetic yields the relational semantics of logic programming (van Emden and Kowalski, 1976); self-reference resolves to multiple targets; and programs are immune to the Expression Problem (Wadler, 1998). This makes inheritance-calculus strictly more expressive than the $\lambda$-calculus in both common sense and Felleisen’s sense (Felleisen, 1991).
declarative programming, inheritance, fixpoint semantics, self-referential records, expression problem, first-order semantics, $\lambda$-calculus, Böhm trees, configuration languages
@@
$c$s: Theory of computation Program semanticsccs: Theory of computation Object oriented constructsccs: Theory of computation Functional constructsccs: Theory of computation Program schemesccs: Software and its engineering Formal language definitionsccs: Software and its engineering Extensible languagesccs: Software and its engineering Object oriented languages
@@ 1. Introduction
Declarative and configuration languages are ubiquitous in modern software engineering. Systems such as NixOS modules (Dolstra et al., 2010; NixOS Contributors, 2024), Jsonnet (Cunningham, 2014), Hydra (Yadan and Droppo, 2023), CUE (van Lohuizen, 2019), Dhall (Gonzalez, 2017), Kustomize (Kubernetes SIG CLI, 2018), and JSON Patch (Bryan et al., 2013) all provide mechanisms for composing structured data through inheritance or merging. Among these, the NixOS module system stands out: its recursive attribute set merging with fixpoint semantics and deferred modules (NixOS Contributors, 2024) achieves considerable expressiveness in practice, and the mechanism has been adopted well beyond NixOS itself: Home Manager (Helgesson, 2017), nix-darwin (Jordan, 2017), disko (Lassulus and nix-community Contributors, 2022), flake-parts (Hensing, 2022), dream2nix (Hauer, 2021), devenv (Kožar, 2022), KubeNix (Xtruder and Hall, 2022), and nixidy (Ingólfsson, 2024) use it to manage user environments, macOS configuration, disk partitioning, flake structure, multi-language packaging, developer environments, and Kubernetes clusters respectively. Yet no computational theory explains why this mechanism is so expressive.
@@
Conventionally, the value domain of configuration languages is assumed to consist of finite, well-founded structures—initial algebras in the sense of universal algebra. We challenge this assumption. Configuration values are better understood as lazily observable, possibly infinite structures, where semantics is determined by the observer and a finite prefix suffices for any finite observation. The nixpkgs package collection (Dolstra, 2006) (over 100,000 packages) behaves as a lazily observed structure in practice, allowing any single package to be evaluated without materializing the entire set. Guided by this observation, we set out to reduce the NixOS module system to a minimal set of primitives. The reduction produced MIXIN, an early implementation in Nix, and subsequently MIXINv2 (Yang, 2026), now implemented in Python: a declarative programming language with only three constructs (record literals, properties, and inheritance) and no functions or let-bindings.
@@
In developing real programs in MIXINv2, including boolean logic, natural number arithmetic, and a web application, the linearization problem of mixin-based systems (Bracha and Cook, 1990; Barrett et al., 1996) did not arise. Inheritance is inherently commutative, idempotent, and associative (Section 3). Self-reference naturally resolves to multiple objects rather than one. Prior systems such as Scala and the NixOS module system assume single-object resolution and reject the multi-path situation. Self-referential records require no fixed-point combinator in the object language. The mixin tree is the call stack, and qualified this resolution is a first-order set-theoretic query over paths. We also found that boolean logic and arithmetic, implemented as separate MIXINv2 files, extend in both dimensions of the Expression Problem (Wadler, 1998) without any dedicated mechanism. Church-encoded values automatically form trie structures, and operations on tries automatically become Cartesian products, so that adding $\{1,2\}$ to $\{3,4\}$ yields $\{4,5,6\}$—the relational semantics of logic programming (van Emden and Kowalski, 1976) (Section 5). Furthermore, the same web application MIXINv2 file runs under both Python asyncio and synchronous execution without modification, an instance of function color blindness (Nystrom, 2015) (Section Function color blindness: ).
@@
This paper distills these observations into inheritance-calculus, a minimal computational model for declarative programming. Function application translates directly into inheritance: a $\lambda$-abstraction corresponds to a record with an $\mathrm{argument}$ slot and a $\mathrm{result}$ slot, and function application corresponds to inheriting from the function record and a record that supplies the argument. Since the $\lambda$-calculus embeds into inheritance-calculus, inheritance-calculus is Turing complete. The translation reveals an asymmetry: embedding the $\lambda$-calculus into inheritance-calculus requires five local, syntax-directed rules (Section 4), but the lazy $\lambda$-calculus cannot macro-express the inheritance constructors. Inheritance-calculus enables open extension that changes operational equivalence on $\lambda$-terms, establishing that inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2).
@@
All constructions in this paper have been implemented and tested in MIXINv2, available as open source (Yang, 2026), with a test suite covering every example. [^1] MIXINv2 separates symbol resolution (which resolves lexical references and qualified-this references) from runtime mixin evaluation, but symbol resolution is performed just-in-time per symbol rather than ahead of time, so it does not provide static type checking in the traditional sense. Nevertheless, MIXINv2’s resolution phase is stricter than the untyped inheritance-calculus. All examples in this paper are ported from MIXINv2 and conform to both semantics. Elements added to satisfy the resolver are harmless for the algorithms presented here and may aid readability. All tests are mechanically verified against expected output snapshots.
@@
  • We present inheritance-calculus, a minimal computational model for declarative programming that contains three constructs (record, definition, inheritance).
  • The mixin is a single abstraction that unifies class, method, object, and property. Method-level overriding is subsumed by deep merge of mixin trees; inheritance is commutative, idempotent, and associative, and the linearization problem in OOP does not arise.
  • The semantics is observational: the observer drives computation by querying paths in a lazily constructed tree. The six semantic equations are first-order [^2] First-order in three senses simultaneously: the semantic domain contains only paths (finite sequences of labels) and sets of paths, with no function spaces, closures, or strategies; the six equations are first-order functions whose inputs and outputs are paths and sets of paths, never functions; and the judgments are first-order inductive definitions (Aczel, 1977) over paths and labels, where quantifiers range over paths and labels, never over functions or predicates., denotational, and fixpoint-computable.
  • The mixin tree is the call stack. Self-reference resolves to multiple objects. Since inheritance is idempotent, all objects contribute equally.
  • The $\lambda$-calculus macro-expresses into inheritance-calculus in five local rules (Section 4).
  • The translation gives the lazy $\lambda$-calculus a first-order semantics (Section 4.1): the prior higher-order semantic frameworks are shown to be faithful to a sublanguage of inheritance-calculus’s first-order equations. Appendix E reformulates this semantics directly on ANF. As a corollary, some $\lambda$-terms that diverge in the $\lambda$-calculus converge under $\mathrm{lfp}(T_{P})$ after translation (Section 4; Remark B.5).
  • Inheritance-calculus is strictly more expressive than the lazy $\lambda$-calculus in Felleisen’s sense (Felleisen, 1991) (Section 4.2).
  • We observe further properties: immunity to nonextensibility in the sense of the Expression Problem (Wadler, 1998) (Section 5), relational semantics of logic programming from ordinary arithmetic (Section 5), and function color blindness (Nystrom, 2015) (Section Function color blindness: ).
@@ 2. Syntax
Let $e$ denote an expression. Let $\ell$ denote a label (property name). Let $k$ denote a nonnegative integer (path length).
@@
The $\mapsto$ in $\ell\mapsto e$ defines a property. MIXINv2 also provides scalar types through a foreign-function interface (FFI); inheritance-calculus is what remains after removing the FFI. None of the examples in this paper uses the FFI; scalars did not arise.
@@ 3. Mixin Trees
In object-oriented programming, a mixin (Bracha and Cook, 1990) is a fragment of behavior that can be composed onto a class via inheritance. Inheritance-calculus has a single abstraction, the deep-mergeable mixin, [^3] Where no ambiguity arises, we write simply “mixin.” The distinction from the Bracha–Cook mixin (Bracha and Cook, 1990) is detailed in Section 7. which can simultaneously serve as field, method, class, and object, because there are no opaque methods: every value is a transparent record, and inheritance always performs a deep merge of subtrees. We call the resulting structure a mixin tree. Its semantics is purely observational: given an inheritance-calculus program, one queries whether a label $\ell$ is a property at a given position $p$.
This section defines the semantics via six mutually recursive first-order functions on paths. The six functions map object-oriented concepts to set-theoretic operations: $\mathtt{properties}$ and $\mathtt{supers}$ generalize member lookup and inheritance chains, $\mathtt{overrides}$ and $\mathtt{bases}$ capture method override and base classes, and $\mathtt{resolve}$ and $\mathtt{this}$ handle name resolution and self-reference.
3.1. Definitions
@@
A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ell_{2},\ldots,\ell_{n})$ that identifies a position in the tree. We write $p$ for a path. The root path is the empty sequence $()$. Given a path $p$ and a label $\ell$, $p\mathbin{\triangleright}\ell$ is the sequence $p$ extended with $\ell$. For a nonroot path, the parent path $\mathtt{init}(p)$ is $p$ with its last element removed, and the final label $\mathtt{last}(p)$ is the last element of $p$. For convenience, we sometimes write $\ell$ for $\mathtt{last}(p)$ when the path is clear from context. With paths in hand, we can state what the AST provides at each path.
@@
Parsing populates $\mathtt{defines}$ and $\mathtt{inherits}$ from each element of a record. A definition $\ell\mapsto e$ contributes $\ell$ to $\mathtt{defines}(p)$. A reference contributes $(n,\;\ell_{\mathrm{down},*})$ to $\mathtt{inherits}(p)$. Both $\mathtt{defines}$ and $\mathtt{inherits}$ are pure data; they are not functions of runtime state. Given these two primitives, we can now answer the question posed at the beginning of this section.
@@
Since a mixin can play the role of a method, a child scope should be able to find it through the parent’s label even if the child does not redefine that label; this is method override. Concretely, inheritance can introduce multiple definitions of the same label at the same scope level; $\mathtt{overrides}(p)$ collects all such paths that share the same identity as $p$, so that their definitions merge rather than shadow each other, enabling deep merge. The overrides of $p$ include $p$ itself and $p_{\mathrm{branch}}\mathbin{\triangleright}\mathtt{last}(p)$ for every branch $p_{\mathrm{branch}}$ of $\mathtt{init}(p)$ that also defines $\mathtt{last}(p)$:
@@
At each step $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a nonnegative integer, the recursion terminates. [^4] If no matching pair $(p_{\mathrm{site}},\;p_{\mathrm{override}})$ exists at some step, the frontier becomes the empty set, and $\mathtt{resolve}$ returns $\varnothing$: the reference has no target, and no properties are inherited. This is the most self-consistent semantics for the untyped calculus. MIXINv2 (Yang, 2026) is stricter: its resolver rejects references whose frontier becomes empty, reporting a resolution error at compile time.
@@
3.2. Recursive Evaluation
The six equations (1)–(6) define $\mathtt{properties}(p)$ by mutual recursion: $\ell\in\mathtt{properties}(p)$ holds if and only if it can be established by a finite chain of applications of these functions. The functions are the interpreter—the observer drives the computation by choosing which path to inspect, and each equation unfolds on demand. Because the functions are pure and order-independent, the reflexive-transitive closure $\mathtt{bases}^{*}$ in (2) may be explored breadth-first or depth-first without affecting the result; paths may be interned so that structural equality reduces to pointer equality; and the six functions form a memoizable dynamic-programming recurrence keyed by paths.
Operationally, evaluation uses tabled evaluation (Tamaki and Sato, 1986; Chen and Warren, 1996): each query is memoised on first encounter, and re-entrant calls (cycles in the dependency graph) return the current accumulator rather than diverging. Denotationally, the six equations define a monotone immediate consequence operator $T_{P}$ (van Emden and Kowalski, 1976) whose right-hand sides use only $\cup$, $\in$, equality, and positive set comprehension, with no negation, set difference, or complement. By the Knaster–Tarski theorem (Tarski, 1955), $T_{P}$ has a unique least fixed point $\mathrm{lfp}(T_{P})$, which serves as the denotational semantics. Appendix A proves that tabled evaluation never returns a wrong answer and analyses termination.